A 10 pound homogeneous ball (IG= 2mr2/5) has a radius of 0.3 feet. It starts from rest and rolls down a 45-degree slope without sliding. Determine the linear speed of the ball in ft/s after rolling 6 feet along the slope.
Answer #1

uzo -6ft R = 0.3ft )مد m = maldi lo lb IR=0.3 ft 10=45° 1 L= U= initial velocity of ball zo v= final 6 ft Applying Energy Convề [s + 7 to x 9.8) 4 G A Singso = 10*9.81*64 lint 50 V 10x9.8) 6* Sin 45° 7 V= 7.7/088 ft/sec And wer